OpenStudy (anonymous):

The exact value of cos 5pi/ 12 is A. sqrt2 (sqrt(3) + 1) / 4 b. sqrt2 (sqrt(3) - 1) / 4 c. .26 d. .99

OpenStudy (raden):

5pi/12 = 75 degrees, right? cos75 = cos(45+30) use the formula : cos(a+b) = cosacosb - sinasinb

OpenStudy (anonymous):

uh..what?

OpenStudy (anonymous):

should it be th edecimel or the other two ?

OpenStudy (raden):

just plug a=45 and b=30 into that formula, next simplify it

OpenStudy (anonymous):

i got .25888 which would have been c....but its not ...its b ...ty

OpenStudy (anonymous):

im confused.

OpenStudy (anonymous):

anyways moving on ..next question

OpenStudy (anonymous):

ty for your time.

OpenStudy (raden):

cos75 = cos(45+30) = cos45cos30 - sin45sin30 = (1/2)sqrt(2)*(1/2)sqrt(3) - (1/2)sqrt(2)*(1/2) = (1/4)sqrt(6) - (1/4)sqrt(2) = (sqrt(6) - sqrt(2))/4 = sqrt(2){sqrt(3) - 1} / 4 B